# Predicting Polymer Properties Based on Multimodal Multitask Pretraining

### v1

@@ 3.3 Learning Paradigm
3.3.1 Masked Prediction Task
During the masked prediction task (i.e., masked language modeling task in NLP), approximately 15% of tokens within a polymer sequence are randomly selected for masking. Subsequently, these chosen tokens are subjected to three possible replacement options: they may be replaced with a special token [MASK], a random token, or left unchanged. Then, the 1D representation network $f^{1d}$ is trained to recover the original identity of these masked tokens based on the contextual information provided by the surrounding sequence. This task encourages our 1D representation network $f^{1d}$ to capture the nuanced structures and interactions present in polymer sequences.
Specifically, $\mathcal{V}$ is the vocabulary set, $\mathcal{M}$ is the set of masked positions, $\bm{\hat{Y}}=[\bm{\hat{y}}_{i}]\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{V}|}$ is the predicted probability distribution over these masked positions, and $\bm{Y}=[\bm{{y}}_{i}]\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{V}|}$ represents the corresponding label in the one-hot format. Here, we utilize the cross-entropy loss [57] as our loss function for the masked prediction task, i.e.,
$\mathcal{L}_{\text{1D}}=-\frac{1}{|\mathcal{M}|}\sideset{}{{}_{i\in\mathcal{M}}}{\sum}\sideset{}{{}_{j}^{|\mathcal{V}|}}{\sum}\bm{y_{i}}[j]\cdot\log(\bm{\hat{y}_{i}}[j]),$ (12)
@@ 3.3.1 Masked Prediction Task
3.3.2 Coordinate Denoising Task
@@
As described Sec. 3.2.2, the corrupted 3D conformation $\bm{\tilde{C}}$ is first converted to initial atom-level representation $\tilde{\bm{X}}^{(0)}_{a}\in\mathbb{R}^{N\times d_{a}}$ and pair-level representation $\tilde{\bm{X}}^{(0)}_{p}\in\mathbb{R}^{N\times N\times d_{p}}$. Then $\tilde{\bm{X}}^{(0)}_{a}$ and $\tilde{\bm{X}}^{(0)}_{p}$ are fed into our attention-based backbone to get the final atom-level representation $\tilde{\bm{X}}^{(L)}_{a}$ and pair-level representation $\tilde{\bm{X}}^{(L)}_{p}$ through $L$ encoder layer. Finally, a pair-level decoder is used to recover the original 3D atom coordinates based on the learned pair-level representation. Here, we utilize the smooth L1 loss [47] as our loss function for the coordinate denoising task, i.e.,
$\bm{\hat{p}}_{i}=\tilde{\bm{p}}_{i}+\sideset{}{{}_{j=1}^{N}}{\sum}\frac{\psi(\tilde{\bm{x}}^{(L)}_{p,ij}-\tilde{\bm{x}}^{(0)}_{p,ij})(\tilde{\bm{p}}_{i}-\tilde{\bm{p}}_{j})}{N},$ (13)
@@
Generally, the coordinate denoising task enables our 3D representation network $f^{3d}$ to accurately capture essential polymer 3D structural information, essential for downstream polymer property prediction tasks.
3.3.3 Cross-modal Alignment Task
In the pretraining phase, in addition to predicting masked tokens and recovering clear 3D coordinates, we also employ constructive learning to align 1D sequential representation $\bm{X}_{1d}$ learned by 1D representation network $f^{1d}$ and 3D structural representation $\bm{X}_{3d}$ learned by 3D representation network $f^{3d}$.
Specifically, we derive the corresponding multimodal representations $\{\bm{X}_{1d}^{i},\bm{X}_{3d}^{i}\}_{i=1}^{K}$ for a batch of $K$ polymers. Our goal is to maximize the similarity between positive multimodal representation pairs from the same polymer and minimize the similarity between negative multimodal representation pairs from different polymers. Here, we utilize the classical InfoNCE loss [29] as our loss function for the cross-modal alignment task, i.e.,
$\mathcal{L}_{\text{contrastive}}=-\frac{1}{K}\sideset{}{{}_{i=1}^{K}}{\sum}\log\Biggl{(}\frac{\exp(\text{sim}(\bm{X}_{1d}^{i},\bm{X}_{3d}^{i})/\tau)}{\sum_{j=1}^{K}\exp(\text{sim}(\bm{X}_{1d}^{i},\bm{X}_{3d}^{j})/\tau)}\Biggr{)},$ (14)
@@ 3.3.4 Fine-tuning
[…]
4 Experiments
@@
Table 1: The summary of our datasets, where the pretraining dataset is unlabeled and the fine-tuning datasets are regression-type datasets with corresponding property labels.
@@
4.1 Experimental Setup
4.1.1 Dataset
We use the PI1M dataset [27], which contains about one million unlabeled polymer data, to pretrain our MMPolymer. Then, we employ eight open-source polymer property datasets (denoted as Egc, Egb, Eea, Ei, Xc, EPS, Nc, and Eat, respectively) provided in [56] as our fine-tuning datasets. These property datasets, obtained through density functional theory (DFT) calculations [30], encompass a broad range of typical polymer properties. Besides, they are the only publicly available polymer property datasets [^2] Although some other polymer property datasets have been mentioned in previous works [19, 20], these datasets like PolyInfo database [31], are not publicly available, thus making them inaccessible for our experiments. to date and have been extensively utilized in previous works [56, 53]. More details are presented in Table 1.
Table 2: The performance comparison of different methods on eight polymer property datasets, and the best result for each polymer property dataset has been bolded.
@@ 4.1.1 Dataset
RMSE ($\downarrow$) ChemBERTa [1] 0.539±${}_{\pm\text{0.049}}$ 0.664±${}_{\pm\text{0.079}}$ 0.350±${}_{\pm\text{0.036}}$ 0.485±${}_{\pm\text{0.086}}$ 18.711±${}_{\pm\text{1.396}}$ 0.603±${}_{\pm\text{0.083}}$ 0.140±${}_{\pm\text{0.010}}$ 0.219±${}_{\pm\text{0.056}}$
MolCLR [50] 0.587±${}_{\pm\text{0.024}}$ 0.644±${}_{\pm\text{0.072}}$ 0.404±${}_{\pm\text{0.017}}$ 0.533±${}_{\pm\text{0.053}}$ 21.719±${}_{\pm\text{1.631}}$ 0.631±${}_{\pm\text{0.045}}$ 0.117±${}_{\pm\text{0.015}}$ 0.094±${}_{\pm\text{0.033}}$
3D Infomax [41] 0.494±${}_{\pm\text{0.039}}$ 0.553±${}_{\pm\text{0.032}}$ 0.335±${}_{\pm\text{0.055}}$ 0.449±${}_{\pm\text{0.086}}$ 19.483±${}_{\pm\text{2.491}}$ 0.582±${}_{\pm\text{0.054}}$ 0.101±${}_{\pm\text{0.018}}$ 0.094±${}_{\pm\text{0.039}}$
Uni-Mol [59] 0.489±${}_{\pm\text{0.028}}$ 0.531±${}_{\pm\text{0.055}}$ 0.332±${}_{\pm\text{0.027}}$ 0.407±${}_{\pm\text{0.080}}$ 17.414±${}_{\pm\text{1.581}}$ 0.536±${}_{\pm\text{0.053}}$ 0.095±${}_{\pm\text{0.013}}$ 0.084±${}_{\pm\text{0.034}}$
SML [56] 0.489±${}_{\pm\text{0.056}}$ 0.547±${}_{\pm\text{0.110}}$ 0.313±${}_{\pm\text{0.016}}$ 0.432±${}_{\pm\text{0.060}}$ 18.981±${}_{\pm\text{1.258}}$ 0.576±${}_{\pm\text{0.020}}$ 0.102±${}_{\pm\text{0.010}}$ 0.062±${}_{\pm\text{0.014}}$
PLM [56] 0.459±${}_{\pm\text{0.036}}$ 0.528±${}_{\pm\text{0.081}}$ 0.322±${}_{\pm\text{0.037}}$ 0.444±${}_{\pm\text{0.062}}$ 19.181±${}_{\pm\text{1.308}}$ 0.576±${}_{\pm\text{0.060}}$ ${}_{\pm\text{0.010}}$ 0.050±${}_{\pm\text{0.010}}$
polyBERT [20] 0.553±${}_{\pm\text{0.011}}$ 0.759±${}_{\pm\text{0.042}}$ 0.363±${}_{\pm\text{0.037}}$ 0.526±${}_{\pm\text{0.068}}$ 18.437±${}_{\pm\text{0.560}}$ 0.618±${}_{\pm\text{0.049}}$ 0.113±${}_{\pm\text{0.003}}$ 0.172±${}_{\pm\text{0.016}}$
Transpolymer [53] 0.453±${}_{\pm\text{0.007}}$ 0.576±${}_{\pm\text{0.021}}$ 0.326±${}_{\pm\text{0.040}}$ 0.397±${}_{\pm\text{0.061}}$ 17.740±${}_{\pm\text{0.732}}$ 0.547±${}_{\pm\text{0.051}}$ 0.096±${}_{\pm\text{0.016}}$ 0.147±${}_{\pm\text{0.093}}$
@@
$R^{2}$ ($\uparrow$) ChemBERTa [1] 0.880±${}_{\pm\text{0.023}}$ 0.881±${}_{\pm\text{0.028}}$ 0.888±${}_{\pm\text{0.035}}$ 0.745±${}_{\pm\text{0.102}}$ 0.365±${}_{\pm\text{0.098}}$ 0.682±${}_{\pm\text{0.123}}$ 0.643±${}_{\pm\text{0.076}}$ 0.590±${}_{\pm\text{0.078}}$
MolCLR [50] 0.858±${}_{\pm\text{0.010}}$ 0.882±${}_{\pm\text{0.027}}$ 0.854±${}_{\pm\text{0.038}}$ 0.689±${}_{\pm\text{0.037}}$ 0.176±${}_{\pm\text{0.026}}$ 0.683±${}_{\pm\text{0.020}}$ 0.764±${}_{\pm\text{0.037}}$ 0.885±${}_{\pm\text{0.104}}$
3D Infomax [41] 0.900±${}_{\pm\text{0.016}}$ 0.898±${}_{\pm\text{0.018}}$ 0.891±${}_{\pm\text{0.045}}$ 0.766±${}_{\pm\text{0.086}}$ 0.274±${}_{\pm\text{0.122}}$ 0.690±${}_{\pm\text{0.063}}$ 0.797±${}_{\pm\text{0.086}}$ 0.869±${}_{\pm\text{0.097}}$
Uni-Mol [59] 0.901±${}_{\pm\text{0.013}}$ 0.925±${}_{\pm\text{0.011}}$ 0.901±${}_{\pm\text{0.027}}$ 0.820±${}_{\pm\text{0.075}}$ 0.454±${}_{\pm\text{0.079}}$ 0.751±${}_{\pm\text{0.085}}$ 0.828±${}_{\pm\text{0.072}}$ 0.937±${}_{\pm\text{0.032}}$
SML [56] 0.901±${}_{\pm\text{0.022}}$ ${}_{\pm\text{0.029}}$ 0.915±${}_{\pm\text{0.015}}$ 0.802±${}_{\pm\text{0.051}}$ 0.340±${}_{\pm\text{0.125}}$ 0.726±${}_{\pm\text{0.038}}$ 0.812±${}_{\pm\text{0.058}}$ 0.967±${}_{\pm\text{0.015}}$
PLM [56] 0.911±${}_{\pm\text{0.014}}$ 0.925±${}_{\pm\text{0.021}}$ ${}_{\pm\text{0.019}}$ 0.791±${}_{\pm\text{0.049}}$ 0.330±${}_{\pm\text{0.105}}$ 0.726±${}_{\pm\text{0.058}}$ 0.817±${}_{\pm\text{0.056}}$ 0.980±${}_{\pm\text{0.008}}$
polyBERT [20] 0.875±${}_{\pm\text{0.006}}$ 0.844±${}_{\pm\text{0.034}}$ 0.880±${}_{\pm\text{0.035}}$ 0.705±${}_{\pm\text{0.085}}$ 0.384±${}_{\pm\text{0.066}}$ 0.681±${}_{\pm\text{0.058}}$ 0.769±${}_{\pm\text{0.034}}$ 0.672±${}_{\pm\text{0.119}}$
Transpolymer [53] 0.916±${}_{\pm\text{0.002}}$ 0.911±${}_{\pm\text{0.008}}$ 0.902±${}_{\pm\text{0.036}}$ 0.830±${}_{\pm\text{0.059}}$ 0.430±${}_{\pm\text{0.058}}$ 0.744±${}_{\pm\text{0.075}}$ 0.826±${}_{\pm\text{0.071}}$ 0.800±${}_{\pm\text{0.172}}$
@@
Figure 5: t-SNE visualization of MMPolymer on eight polymer property datasets, where the color is determined by the corresponding ground truth.
4.1.2 Baselines
To demonstrate the effectiveness of our method, several state-of-the-art methods have been compared, including SML [56], PLM [56], polyBERT [20] and Transpolymer [53], which are all pretraining-based methods for polymer property prediction. Methods lacking pretraining, like GP [19], have already been surpassed by current pretraining-based methods [53], hence we no longer include them in our comparisons. Besides, we also compare our method with four representative molecular pretraining methods, including ChemBERTa [1],  MolCLR [50],  3D Infomax [41] and  Uni-Mol [59], to reveal the differences between small molecules and polymers. Besides, these baselines are implemented based on their default settings in the corresponding references.
4.1.3 Evaluation Metircs
Since polymer property prediction tasks on the fine-tuning datasets are all regression tasks, we also choose root mean squared error (RMSE) and R-squared ($R^{2}$) as evaluation metrics based on the 5-fold cross-validation in line with previous works [53, 20, 56], thus guaranteeing a comprehensive and fair assessment of the predictive performance.
4.1.4 Model architecture and hyperparameter settings
Our 1D representation network consists of 6 encoding layers, with 12 attention heads each, whereas our 3D representation network comprises 15 encoding layers, with 64 attention heads each. Besides, our model is pretrained on the eight Tesla V100 GPUs using a batch size of 16 and further fine-tuned on the single Tesla V100 GPU using a batch size of 32. Meanwhile, the adam optimizer [18] is utilized with a learning rate of 1e-4, betas of $(0.9,0.99)$, and eps of 1e-6.
4.2 Comparisons
4.2.1 Demonstration of effectiveness
Table 2 presents the performance of different methods on the eight polymer property datasets, which are evaluated by RMSE and $R^{2}$. In general, our proposed method, MMPolymer, stands out by achieving state-of-the-art performance on seven datasets and comparable performance on the remaining Eat dataset, underscoring its effectiveness on polymer property prediction tasks. Notably, several previous works, such as Transpolymer [53], have excluded the Eat dataset from their original paper because it fails to capture the complete polymer structure [^3] Eat is a localized property, which depends on the local atomic environment. and therefore leads to lower reliability compared to other property datasets. Besides, even if considering the Eat dataset, the superiority of our MMPolymer is still obvious.
Moreover, the comparison with the four representative molecular pretraining methods reveals insightful trends. On the one hand, the 3D molecular pretraining method Uni-Mol consistently outperforms the 1D molecular pretraining method ChemBERTa across all eight polymer property datasets, emphasizing the crucial role of 3D structural information in enhancing predictive performance. On the other hand, all molecular pretraining methods exhibit inferior performance when compared with our MMPolymer and the best polymer pretraining baseline. These observations indicate the inherent limitations of directly applying molecular pretraining methods to polymer-specific tasks, thereby emphasizing the significance of tailored methods like our MMPolymer for achieving accurate polymer property prediction.
As shown in Figure 5, we conduct t-SNE visualization [43] based on the polymer representations learned by our MMPolymer and corresponding ground truth for each polymer property dataset. The t-SNE visualization illustrates that the chemical space defined by the representations accurately captures the differences in the ground truth for each data point, as reflected by the color gradations. These colors are determined by the corresponding ground truth, and the spread of colors across the datasets indicates a diverse range of values. This variation confirms that the representations learned by our MMPolymer are sensitive to the underlying ground truth differences among data points. The distinct clusters, especially in datasets like Egb, Eea, and Eat, where the color intensity marks out separate regions, serve as a testament to the fidelity with which the representations map out the actual property landscape of the dataset. The linear color variation further emphasizes that the representations are not only capturing distinct property values but are also able to reflect a continuum in the property and chemical space, showing the superiority of our MMPolymer.
Table 3: The performance improvement of MMPolymer-1D, which only utilizes polymer 1D sequential information during fine-tuning, and MMPolymer-3D, which only utilizes 3D structural information during fine-tuning, compared with the best baseline on each polymer property dataset.
@@ 4.2.1 Demonstration of effectiveness
4.2.2 Adaptability for various downstream settings
Table 4: The performance of MMPolymer under different data processing strategies, where the "Star Keep" strategy refers to directly using the original P-SMILES string, the "Star Remove" strategy refers to removing "*" symbols in the original P-SMILES string and the "Star Substitution" strategy refers to replacing the "*" symbol in the original P-SMILES string with the neighboring atom symbol of another "*" symbol.

### v2

@@ 3.3. Learning Paradigm
3.3.1. Masked Prediction Task
During the masked prediction task (i.e., masked language modeling task in NLP), approximately 15% of tokens in a polymer sequence are randomly selected for masking. Subsequently, these chosen tokens are subjected to three possible replacement options: they may be replaced with a special token [MASK], a random token, or left unchanged. Then, the 1D representation network $f^{1d}$ is trained to recover the original identity of these masked tokens based on the contextual information provided by the surrounding sequence. This task encourages our 1D representation network $f^{1d}$ to capture the nuanced structures and interactions in polymer sequences.
Specifically, $\mathcal{V}$ is the vocabulary set, $\mathcal{M}$ is the set of masked positions, $\bm{\hat{Y}}=[\bm{\hat{y}}_{i}]\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{V}|}$ is the predicted probability distribution over these masked positions, and $\bm{Y}=[\bm{{y}}_{i}]\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{V}|}$ represents the corresponding label in the one-hot format. Here, we utilize the cross-entropy loss (Zhang and Sabuncu, 2018) as our loss function for the masked prediction task, i.e.,
(12) $\mathcal{L}_{\text{1D}}=-\frac{1}{|\mathcal{M}|}\sideset{}{{}_{i\in\mathcal{M}}}{\sum}\sideset{}{{}_{j}^{|\mathcal{V}|}}{\sum}\bm{y_{i}}[j]\cdot\log(\bm{\hat{y}_{i}}[j]),$
@@ 3.3.1. Masked Prediction Task
3.3.2. Coordinate Denoising Task
@@
As described Sec. 3.2.2, the corrupted 3D conformation $\bm{\tilde{C}}$ is first converted to initial atom-level representation $\tilde{\bm{X}}^{(0)}_{a}\in\mathbb{R}^{N\times d_{a}}$ and pair-level representation $\tilde{\bm{X}}^{(0)}_{p}\in\mathbb{R}^{N\times N\times d_{p}}$. Then $\tilde{\bm{X}}^{(0)}_{a}$ and $\tilde{\bm{X}}^{(0)}_{p}$ are fed into our attention-based backbone to get the final atom-level representation $\tilde{\bm{X}}^{(L)}_{a}$ and pair-level representation $\tilde{\bm{X}}^{(L)}_{p}$ through $L$ encoder layer. Finally, we use a pair-level decoder to recover the original 3D atom coordinates based on the learned pair-level representation. Here, we utilize the smooth L1 loss (Wang et al., 2020) as our loss function for the coordinate denoising task, i.e.,
(13) $\bm{\hat{p}}_{i}=\tilde{\bm{p}}_{i}+\sideset{}{{}_{j=1}^{N}}{\sum}\frac{\psi(\tilde{\bm{x}}^{(L)}_{p,ij}-\tilde{\bm{x}}^{(0)}_{p,ij})(\tilde{\bm{p}}_{i}-\tilde{\bm{p}}_{j})}{N},$
@@
Generally, the coordinate denoising task enables our 3D representation network $f^{3d}$ to accurately capture essential polymer 3D structural information, which is crucial for downstream polymer property prediction tasks.
3.3.3. Cross-modal Alignment Task
During pretraining, we also employ contrastive learning to align 1D sequential representation $\bm{X}_{1d}$ learned by 1D representation network $f^{1d}$ and 3D structural representation $\bm{X}_{3d}$ learned by 3D representation network $f^{3d}$.
Specifically, we derive the corresponding multimodal representations $\{\bm{X}_{1d}^{i},\bm{X}_{3d}^{i}\}_{i=1}^{K}$ for a batch of $K$ polymers. Our goal is to maximize the similarity between positive multimodal representation pairs from the same polymer and minimize the similarity between negative multimodal representation pairs from different polymers. Here, we utilize the classical InfoNCE loss (Oord et al., 2018), as our loss function for the cross-modal alignment task, i.e.,
(14) $\mathcal{L}_{\text{contrastive}}=-\frac{1}{K}\sideset{}{{}_{i=1}^{K}}{\sum}\log\Biggl{(}\frac{\exp(\text{sim}(\bm{X}_{1d}^{i},\bm{X}_{3d}^{i})/\tau)}{\sum_{j=1}^{K}\exp(\text{sim}(\bm{X}_{1d}^{i},\bm{X}_{3d}^{j})/\tau)}\Biggr{)},$
@@ 3.3.4. Fine-tuning
Moreover, although the pretraining phase is multi-modal, the fine-tuning phase allows for flexible modality information choice by using either a single modality representation (1D sequential representation $\bm{X}_{1d}$ or 3D structural representation $\bm{X}_{3d}$) or combining both modalities to predict corresponding polymer properties. This choice depends on the characteristics of the property being predicted and the specific requirements of the polymer property prediction task at hand. Regardless of the chosen modality representation, the multimodal multitask pretraining paradigm has already equipped our MMPolymer with adequate general knowledge for achieving accurate polymer property prediction.
4. Experiments
@@
Table 1. The summary of our datasets, where the pretraining dataset is unlabeled and the fine-tuning datasets are regression-type datasets with corresponding property labels.
@@
4.1. Experimental Setup
4.1.1. Datasets
We use the PI1M dataset (Ma and Luo, 2020), which contains about one million unlabeled polymer data, to pretrain our MMPolymer. Then, we employ eight open-source polymer property datasets (denoted as Egc, Egb, Eea, Ei, Xc, EPS, Nc, and Eat, respectively) provided in (Zhang et al., 2023) as our fine-tuning datasets [^2] Although some other polymer property datasets have been mentioned in previous works (Kuenneth et al., 2021; Kuenneth and Ramprasad, 2023), these datasets like PolyInfo database (Otsuka et al., 2011), are not publicly available, thus making them inaccessible for our experiments.. These property datasets, obtained through density functional theory (DFT) calculations (Orio et al., 2009), encompass a broad range of typical polymer properties and have been extensively utilized in previous works (Zhang et al., 2023; Xu et al., 2023). More details about our datasets are presented in Table 1.
Table 2. The performance comparison of different methods on eight polymer property datasets, and the best result for each polymer property dataset has been bolded.
@@ 4.1.1. Datasets
RMSE ($\downarrow$) ChemBERTa (Ahmad et al., 2022) 0.539±${}_{\pm\text{0.049}}$ 0.664±${}_{\pm\text{0.079}}$ 0.350±${}_{\pm\text{0.036}}$ 0.485±${}_{\pm\text{0.086}}$ 18.711±${}_{\pm\text{1.396}}$ 0.603±${}_{\pm\text{0.083}}$ 0.140±${}_{\pm\text{0.010}}$ 0.219±${}_{\pm\text{0.056}}$
MolCLR (Wang et al., 2022) 0.587±${}_{\pm\text{0.024}}$ 0.644±${}_{\pm\text{0.072}}$ 0.404±${}_{\pm\text{0.017}}$ 0.533±${}_{\pm\text{0.053}}$ 21.719±${}_{\pm\text{1.631}}$ 0.631±${}_{\pm\text{0.045}}$ 0.117±${}_{\pm\text{0.015}}$ 0.094±${}_{\pm\text{0.033}}$
3D Infomax (Stärk et al., 2022) 0.494±${}_{\pm\text{0.039}}$ 0.553±${}_{\pm\text{0.032}}$ 0.335±${}_{\pm\text{0.055}}$ 0.449±${}_{\pm\text{0.086}}$ 19.483±${}_{\pm\text{2.491}}$ 0.582±${}_{\pm\text{0.054}}$ 0.101±${}_{\pm\text{0.018}}$ 0.094±${}_{\pm\text{0.039}}$
Uni-Mol (Zhou et al., 2023) 0.489±${}_{\pm\text{0.028}}$ 0.531±${}_{\pm\text{0.055}}$ 0.332±${}_{\pm\text{0.027}}$ 0.407±${}_{\pm\text{0.080}}$ 17.414±${}_{\pm\text{1.581}}$ 0.536±${}_{\pm\text{0.053}}$ 0.095±${}_{\pm\text{0.013}}$ 0.084±${}_{\pm\text{0.034}}$
SML (Zhang et al., 2023) 0.489±${}_{\pm\text{0.056}}$ 0.547±${}_{\pm\text{0.110}}$ 0.313±${}_{\pm\text{0.016}}$ 0.432±${}_{\pm\text{0.060}}$ 18.981±${}_{\pm\text{1.258}}$ 0.576±${}_{\pm\text{0.020}}$ 0.102±${}_{\pm\text{0.010}}$ 0.062±${}_{\pm\text{0.014}}$
PLM (Zhang et al., 2023) 0.459±${}_{\pm\text{0.036}}$ 0.528±${}_{\pm\text{0.081}}$ 0.322±${}_{\pm\text{0.037}}$ 0.444±${}_{\pm\text{0.062}}$ 19.181±${}_{\pm\text{1.308}}$ 0.576±${}_{\pm\text{0.060}}$ ${}_{\pm\text{0.010}}$ 0.050±${}_{\pm\text{0.010}}$
polyBERT (Kuenneth and Ramprasad, 2023) 0.553±${}_{\pm\text{0.011}}$ 0.759±${}_{\pm\text{0.042}}$ 0.363±${}_{\pm\text{0.037}}$ 0.526±${}_{\pm\text{0.068}}$ 18.437±${}_{\pm\text{0.560}}$ 0.618±${}_{\pm\text{0.049}}$ 0.113±${}_{\pm\text{0.003}}$ 0.172±${}_{\pm\text{0.016}}$
Transpolymer (Xu et al., 2023) 0.453±${}_{\pm\text{0.007}}$ 0.576±${}_{\pm\text{0.021}}$ 0.326±${}_{\pm\text{0.040}}$ 0.397±${}_{\pm\text{0.061}}$ 17.740±${}_{\pm\text{0.732}}$ 0.547±${}_{\pm\text{0.051}}$ 0.096±${}_{\pm\text{0.016}}$ 0.147±${}_{\pm\text{0.093}}$
@@
$R^{2}$ ($\uparrow$) ChemBERTa (Ahmad et al., 2022) 0.880±${}_{\pm\text{0.023}}$ 0.881±${}_{\pm\text{0.028}}$ 0.888±${}_{\pm\text{0.035}}$ 0.745±${}_{\pm\text{0.102}}$ 0.365±${}_{\pm\text{0.098}}$ 0.682±${}_{\pm\text{0.123}}$ 0.643±${}_{\pm\text{0.076}}$ 0.590±${}_{\pm\text{0.078}}$
MolCLR (Wang et al., 2022) 0.858±${}_{\pm\text{0.010}}$ 0.882±${}_{\pm\text{0.027}}$ 0.854±${}_{\pm\text{0.038}}$ 0.689±${}_{\pm\text{0.037}}$ 0.176±${}_{\pm\text{0.026}}$ 0.683±${}_{\pm\text{0.020}}$ 0.764±${}_{\pm\text{0.037}}$ 0.885±${}_{\pm\text{0.104}}$
3D Infomax (Stärk et al., 2022) 0.900±${}_{\pm\text{0.016}}$ 0.898±${}_{\pm\text{0.018}}$ 0.891±${}_{\pm\text{0.045}}$ 0.766±${}_{\pm\text{0.086}}$ 0.274±${}_{\pm\text{0.122}}$ 0.690±${}_{\pm\text{0.063}}$ 0.797±${}_{\pm\text{0.086}}$ 0.869±${}_{\pm\text{0.097}}$
Uni-Mol (Zhou et al., 2023) 0.901±${}_{\pm\text{0.013}}$ 0.925±${}_{\pm\text{0.011}}$ 0.901±${}_{\pm\text{0.027}}$ 0.820±${}_{\pm\text{0.075}}$ 0.454±${}_{\pm\text{0.079}}$ 0.751±${}_{\pm\text{0.085}}$ 0.828±${}_{\pm\text{0.072}}$ 0.937±${}_{\pm\text{0.032}}$
SML (Zhang et al., 2023) 0.901±${}_{\pm\text{0.022}}$ ${}_{\pm\text{0.029}}$ 0.915±${}_{\pm\text{0.015}}$ 0.802±${}_{\pm\text{0.051}}$ 0.340±${}_{\pm\text{0.125}}$ 0.726±${}_{\pm\text{0.038}}$ 0.812±${}_{\pm\text{0.058}}$ 0.967±${}_{\pm\text{0.015}}$
PLM (Zhang et al., 2023) 0.911±${}_{\pm\text{0.014}}$ 0.925±${}_{\pm\text{0.021}}$ ${}_{\pm\text{0.019}}$ 0.791±${}_{\pm\text{0.049}}$ 0.330±${}_{\pm\text{0.105}}$ 0.726±${}_{\pm\text{0.058}}$ 0.817±${}_{\pm\text{0.056}}$ 0.980±${}_{\pm\text{0.008}}$
polyBERT (Kuenneth and Ramprasad, 2023) 0.875±${}_{\pm\text{0.006}}$ 0.844±${}_{\pm\text{0.034}}$ 0.880±${}_{\pm\text{0.035}}$ 0.705±${}_{\pm\text{0.085}}$ 0.384±${}_{\pm\text{0.066}}$ 0.681±${}_{\pm\text{0.058}}$ 0.769±${}_{\pm\text{0.034}}$ 0.672±${}_{\pm\text{0.119}}$
Transpolymer (Xu et al., 2023) 0.916±${}_{\pm\text{0.002}}$ 0.911±${}_{\pm\text{0.008}}$ 0.902±${}_{\pm\text{0.036}}$ 0.830±${}_{\pm\text{0.059}}$ 0.430±${}_{\pm\text{0.058}}$ 0.744±${}_{\pm\text{0.075}}$ 0.826±${}_{\pm\text{0.071}}$ 0.800±${}_{\pm\text{0.172}}$
@@
Figure 4. t-SNE visualization of MMPolymer on eight polymer property datasets, where the color for each data point is determined by the corresponding ground truth (i.e., property label).
Table 3. The performance improvement of MMPolymer-1D, which only utilizes polymer 1D sequential information during fine-tuning, and MMPolymer-3D, which only utilizes 3D structural information during fine-tuning, compared with the best baseline on each polymer property dataset.
@@
4.1.2. Baselines
To demonstrate the effectiveness of our method, several state-of-the-art methods have been compared, including SML (Zhang et al., 2023), PLM (Zhang et al., 2023), polyBERT (Kuenneth and Ramprasad, 2023) and Transpolymer (Xu et al., 2023), which are all pretraining-based methods for polymer property prediction. Methods lacking pretraining, like GP (Kuenneth et al., 2021), have already been surpassed by current pretraining-based methods (Xu et al., 2023), hence we no longer include them in our comparisons. Besides, we also compare our method with four representative molecular pretraining methods, including ChemBERTa (Ahmad et al., 2022),  MolCLR (Wang et al., 2022),  3D Infomax (Stärk et al., 2022) and  Uni-Mol (Zhou et al., 2023), to reveal the differences between small molecules and polymers. The baselines are implemented based on their default settings in the corresponding references.
4.1.3. Evaluation Metircs
Since polymer property prediction tasks on the fine-tuning datasets are all regression tasks, we also choose root mean squared error (RMSE) and R-squared ($R^{2}$) as evaluation metrics in line with previous works (Xu et al., 2023; Kuenneth and Ramprasad, 2023; Zhang et al., 2023), thus guaranteeing a comprehensive and fair assessment of the predictive performance.
4.1.4. Model architecture and hyperparameter settings
Our 1D representation network consists of 6 encoding layers, with 12 attention heads each, whereas our 3D representation network comprises 15 encoding layers, with 64 attention heads each. Besides, our model is pretrained on the eight Tesla V100 GPUs using a batch size of 16 while the adam optimizer (Kingma and Ba, 2014) is utilized with a learning rate of 1e-4, betas of $(0.9,0.99)$, and eps of 1e-6.
Table 4. The performance of MMPolymer under different data processing strategies, where the "Star Keep" strategy refers to directly using the original P-SMILES string, the "Star Remove" strategy refers to removing "*" symbols in the original P-SMILES string and the "Star Substitution" strategy refers to replacing the "*" symbol in the original P-SMILES string with the neighboring atom symbol of another "*" symbol.